The management of the Diners Delight restaurant chain is in the process of establishing quality control charts for the time that its service people give to each customer. Management thinks the length of time that each customer is given should remain within certain limits to enhance service quality. For simplicity, the data from three samples are displayed below. In each sample, four service people were randomly selected, and the customer service they provided was observed. The time (in seconds) to service each customer was recorded as below.

Service person Sample 1 (morning) Sample 2 (afternoon) Sample 3 (evening)
#1 201 196 197
#2 198 206 210
#3 190 207 211
#4 195 202 206
  • Keep two decimals if not exact, do not round. For example, 3.24923... will be kept as 3.24, but the exact value of 0.625 will be kept as 0.625
  1. What is the sample size of this study?
  2. What is the X-bar value of sample 1?
  3. What is the R-value of sample 2?
  4. What is 99.73% (3-sigma) the upper control limit of the X-bar chart  and R-chart  ?
  5. Assume the established control chart are as follows: X-bar chart is [120, 170], and R-chart is [20, 80]. A sample of six service personnel was observed, and the following customer service times in seconds were recorded: 180, 125, 140, 175, 156 and 160. What is the sample mean here?  Based on the information of sample mean and sample range, and the assumed established control chart, please determine whether the sample process is in control
Answer #1

Sample size of study is 4.

Sample / Person 1 2 3 4 Mean Range
1 201 198 190 195 196 11
2 196 206 207 202 202.75 11
3 197 210 211 206 206 14

xbar value of sample 1 = 196

R value for sample 2 =11

Grand mean xbarbar =201.58

Mean Range Rbar=12

UCLxbar = xbarbxr + A2*Rbar = 201.58 +0.729*12 =210.328

UCLR = D4*Rbar - 2.282*12 =27.384

(ii) Mean in this case =156

Since two values are beyond control limits, the process is out of control.